# Adaptive operator learning for infinite-dimensional Bayesian inverse problems

### v2

@@ 3.1 Adaptive model error reduction
where $\gamma^{j}=\gamma^{j-1}\cup\{\widetilde{m}_{j}\}$, $d(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}$ is the distance function, and $D\big{(}\widehat{\mathcal{G}}_{t}(m),\widehat{\mathcal{G}}_{t}(\gamma^{j-1})\big{)}=\max_{i\leq j-1}d\left(\widehat{\mathcal{G}}_{t}(m),\hat{\mathcal{G}}_{t}(\tilde{m}_{i})\right)$ is the distance between the point $\widehat{\mathcal{G}}_{t}(m)$ and the set $\{\widehat{\mathcal{G}}_{t}(\widetilde{m}_{i})\}_{i=1}^{j-1}$. We can improve the DeepOnet model $\mathcal{F}_{\theta}$ after collecting the new training set $\gamma^{Q}$. More precisely, during the online training phase, we initialize the neural network’s parameter from the previously trained model $\mathcal{F}_{\theta}$. We expect a significant speedup in solving Eq.(5) with this initialization, which can be viewed as an example of transfer learning.
The advantages of this method are obvious. Specifically, the selected adaptive points will have varying features in the surrogate solution spaces and will be close to the anchor point $r^{t}$. This is more of an adversarial process; in order to guarantee local accuracy, we want the new points to be close to the $r^{t}$. In order to enhance generalization capacity, we also want them to incorporate as much data as the current surrogate did. The only remaining question is how to approximate the approximate posterior distribution $\widehat{\nu}_{t}$. To this end, traditional MCMC-based sampling methods and particle-based approaches[41, 42, 43, 44, 45] can be applied. In this work, we only investigate the unscented Kalman inversion (UKI) method[35] because it never needs early stopping or empirical variance inflation. Notably, our framework is easily extensible to numerous variants of popular particle-based methods[46, 47, 48, 49, 43, 50, 51, 52, 53]; however, this is outside the scope of this work.
@@ 3.2 Unscented Kalman Inversion
Let $Y_{n}:=\{y_{1},y_{2},\cdots,y_{n}\}$ denotes the observation set at time $n$. In order to approximate the conditional distribution $\nu_{n}$ of $m_{n}|Y_{n}$, the iterative algorithm starts from the prior $\nu_{0}$ and updates $\nu_{n}$ through the prediction and analysis steps: $\nu_{n}\rightarrow\widetilde{\nu}_{n+1}$, and then $\widetilde{\nu}_{n+1}\rightarrow\nu_{n+1}$, where $\widetilde{\nu}_{n+1}$ is the distribution of $m_{n+1}|Y_{n}$. In the prediction step, we assume that $\nu_{n}=\mathcal{N}(r_{n},C_{n})$, then under Eq. (12), $\widetilde{\nu}_{n+1}$ is also Gaussian with mean and covariance:
@@
We obtain the following UKI algorithm in Algorithm 1 by applying the aforementioned quadrature rules. UKI is a derivative-free algorithm that applies a Gaussian approximation theorem iteratively to transport a set of particles to approximate given distributions. As a result, it only needs $12N_{m}+$ 1 model evaluations per iteration, making it simple to implement and inexpensive to compute.
@@ 3.4 Convergence analysis under the linear case
Under Assumptions 3.2-3.4, suppose $Range(\mathcal{G}^{T})=Range(\widehat{\mathcal{G}}^{T})=\mathbb{R}^{N_{m}}$ and $\Sigma_{\omega}\succ 0,\Sigma_{\eta}\succ 0$. Then the sequence $\widehat{r}_{\infty},\widehat{C}_{\infty}^{-1}$ in Eq.(3.4) obtained by using the surrogate model will converge to the $r_{\infty},C_{\infty}^{-1}$ in Eq.(3.4) and we have the following error estimate
@@ 4 Numerical experiments
[…]
where $y_{ref}=\mathcal{G}(m_{ref})$ are the exact data, $\delta$ dictates the relative noise level and $\xi$ is a Gaussian random variable with zero mean and unit standard deviation. In UKI, the regularization parameters are $\alpha=0.5$ for noise levels 0.05 and 0.1 and $\alpha=1$ for noise levels 0.01. The starting vector for the UKI is chosen at random from $\mathcal{N}(0,I)$. The selection of other hyperparameters is based on [35]. For numerical examples, we set $\Omega=[0,1]^{2}$. The maximum number of UKI iterations is 20 for FEM-UKI and DeepOnet-UKI-Direct and is 10 for DeepOnet-UKI-Adaptive in each cycle. After that, we will choose $Q=50$ adaptive samples for noise level 0.01 and $Q=20$ adaptive samples for noise levels 0.05 and 0.1, respectively, from 2000 samples using the greedy algorithm.
@@
to proof that our adaptive framework can actually reduce the local model error. Moreover, we also calculate the data fitting error via Eq.(3) using the true model during the UKI iteration process.
@@ 4.1 Example 1: Darcy flow
$\mathbf{m}(\mathbf{x})=\sum_{k\in\mathbb{Z}^{+}}\theta_{k}\sqrt{\lambda_{k}}\psi_{k}(\mathbf{x}),$ (33)
where $\lambda_{k}$ and $\psi_{k}$ are the eigenvalues and eigenfunctions, and $\theta_{k}\sim\mathcal{N}(0,1)$ are independent random variables. In practice, we truncate the sum (33) to $n_{d}$ terms, based on the largest $n_{d}$ eigenvalues, and hence $\theta\in\mathbb{R}^{n_{d}}$. The forward problem is solved by FEM method on a $70\times 70$ grid.
We will create the observation data for the inverse problem using the in-distribution data (IDD) and out-of-distribution data (OOD), respectively, as shown in Fig.4. The IDD field $\mathbf{m}_{ref}(\mathbf{x})$ is calculated using Eq.(33) with $n_{d}=256$ and $\theta_{k}\sim\mathcal{N}(0,1)$. The OOD field $\mathbf{m}_{ref}(\mathbf{x})$ is generated for convenience by sampling $\theta_{k}\sim\mathcal{U}[-20,20],k=1,\ldots,256$. To avoid the inverse crime, we will try to inverse the first $N_{m}=128$ KL modes using these observation data.
[…]
Fig. 4: The ground truth for in-distribution data(IDD) and out-of-distribution data (OOD) from above to below. Left: the true permeability fields $\mathbf{m}(\mathbf{x})$. Right: the pressure fields $u(\mathbf{x})$ and the corresponding $36$ equidistant observations with noise level 0.01.
@@
Fig. 5: The data fitting error (left), model error (middle) and relative inversion error (right) for Example 1. Above: IDD case. Below: OOD case.
@@
Fig. 6: IDD case. Above: the estimated permeability field obtained by different methods. Below: the absolute errors with respect to the true ones.
@@
Fig. 7: OOD case. Above: the estimated permeability field obtained by different methods. Below: the absolute errors with respect to the true ones.
@@
Fig. 8: The error box plot of $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$ with increasing adaptive points. Left: the relative inversion errors for IDD inversion. Right: the relative inversion errors for the OOD inversion. Current noise level is 0.01.
@@
Fig. 9: Above: the error box plot of $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$ for IDD inversion and OOD inversion. Below: the mean total number of forward evaluations for FEM-UKI and DeepOnet-UKI-Adaptive respectively.
We repeat the experiment ten times for each $Q\in[20,50,100,150]$ to test the impact of the number $Q$ of adaptive samples used in each refinement. The error box of $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$ is plotted in Fig.8. It is evident from the IDD data that the relative inversion error does not decrease significantly with increasing data. This suggests that a small set of adaptive samples– roughly 50 – can meet the requirements for accuracy and efficiency. On the other hand, the relative inversion error for OOD data steadily drops with increasing dataset size.
In order to demonstrate the computational cost and examine the effects of varying noise levels, we are going to perform the experiment with three different noise levels (0.01, 0.05, and 0.1) and then repeat it with ten different UKI initial values. The numerical results are shown in Fig.9. We can clearly observe that the relative inversion error gradually drops as noise levels rise, suggesting that higher noise levels are less sensitive to model errors. Consequently, our framework performs better in real-world applications with higher noise levels. We plot the mean total online forward evaluations in Fig.9 to show the computational cost. It is evident that our method has a very small cost in comparison to conventional numerical methods, even with adaptive refinement. This is because we only need a maximum of 50 samples to refine the model, in contrast to FEM-UKI, which requires 5140 forward evaluations. Compared to the costly online forward simulations, the online fine-tuning cost is negligible because retraining only takes a few seconds each time.
@@ 4.2 Example 2: The heat source inversion problem
where the initial condition is taken as
$\Phi(x,y)=100\sin(x)\sin(y).$ (35)
The objective is to identify the heat source $f$ from noisy measurements. The heat source field is considered in this paper with the formula $e^{-t}\mathbf{m}(\mathbf{x})$. Conversely, the inverse problem involves using noisy measurements of $u(\mathbf{x},1)$ to determine the true spatial source field $\mathbf{m}(\mathbf{x})$. We assume that the Gaussian random field defined in Eq.(33) is the prior of $\mathbf{m}(\mathbf{x})$. The FEM method is used to solve the forward problem on a $70\times 70$ grid, and the resulting differential equations are integrated using the implicit-Euler scheme with a uniform time step of $\Delta t=0.02$.
We assume that the ground truth $\mathbf{m}_{ref}(\mathbf{x})$ has an analytical solution in this example, i.e.,
@@
Fig. 10: Example 2. Left: the true source field. Right: the temperature field $u(\mathbf{x},1)$ and the corresponding $36$ equidistant observations with noise level 0.01.
To test the effectiveness of our framework, we first run the experiment using the original pre-trained model directly running UKI, i.e. DeepOnet-UKI-Direct. We plot the local model error and the relative inversion error in the middle and right displays of Fig.11, respectively. As expected, the local model error will increase significantly, and the pre-trained model will eventually fail to predict the result. Because of the growing model error, the relative inversion error follows accordingly, increasing dramatically and providing a totally inaccurate final estimate. Nonetheless, we may refine our model by constructing adaptive samples based on this estimate. The procedures are similar to Example 1. During the exploration stage, we will run UKI for 10 steps and select the anchor point with the smallest data fitting error as the initial value for the next run of UKI. To reduce model error, we will generate adaptive samples near the anchor point and refine the surrogate. As shown in the middle of Fig.11, the local model error significantly decreases after refinement. As a result, our method produces relative inversion errors that are significantly reduced after refinement and comparable to those of FEM-UKI. Finally, after several refinements, the entire procedure is terminated based on the stop criteria, with the black star indicating the final point that we would accept for the inversion process using our DeepOnet-UKI-Adaptive algorithm. To demonstrate the effectiveness of our method, we plot the inversion results in Figs.12 and 13. We can see that the final numerical results produced by DeepOnet-UKI-Adaptive and FEM-UKI are very similar and do not differ significantly. This suggests that our method can handle OOD data in closed form.
@@
Fig. 11: The inversion results obtained by the three methods for the heat source case. Left: the data fitting error. Middle: the model error. Right: the relative inversion errors.
@@
Fig. 12: The inversion results for the heat source case. Above: the estimated source fields obtained by three different methods. Below: the absolute errors with respect to the true one.
@@
Fig. 13: The inversion results for the heat source case. Above: the approximate states evaluated at the final estimated source fields obtained by different methods. Below: the absolute errors between the approximate state and the true state.
@@
Fig. 14: The inversion results for the heat source case. Left: the error box plot of $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$. Right: the mean total number of forward evaluations for FEM-UKI and DeepOnet-UKI-Adaptive.
In order to provide additional evidence of the efficacy of our approach, we figure out to perform the experiment for UKI at varying noise levels. In addition, we will repeat the experiment ten times with different initial values for each noise level. Following that, we will compare the number of forward evaluations and the relative inversion errors for every approach. We plot the difference of the relative inversion errors $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$ in the left display of Fig.14. It is clear that when noise levels increase, DeepOnet-UKI-Adaptive performs often better than FEM-UKI. This implies that our method can achieve higher accuracy than traditional solvers. For the reasons mentioned below, the computational cost of the new method can also be extremely small. First of all, it is much faster to fine-tune the original pre-trained surrogate model than it is to solve PDEs. Specially, we only need a maximum of 50 online forward evaluations in this example to retrain the network, which drastically lowers computational costs. We are able to clearly see that DeepOnet-UKI-Adaptive has a substantially smaller total number of forward evaluations than FEM-UKI, as the middle display of Fig.14 illustrates. Secondly, the entire process is automatically stopped by applying the stop criterion. We can start with the initial model that has been trained offline and fine-tune it multiple times for a given inversion task. As a result, our framework can achieve an accuracy level comparable to traditional FEM solvers, but at a significant reduction in computational cost for such problems.

### v3

@@ 3.1 Adaptive model error reduction
where $d\big{(}\widehat{\mathcal{G}}_{t}(\cdot),\widehat{\mathcal{G}}_{t}^{j}\big{)}=\max_{\tilde{m}\in\gamma^{j}}\left\|\widehat{\mathcal{G}}_{t}(\cdot)-\hat{\mathcal{G}}_{t}(\tilde{m})\right\|_{2}$ is the distance between the value $\widehat{\mathcal{G}}_{t}(\cdot)$ and the set $\widehat{\mathcal{G}}_{t}^{j}$. Here $\lambda$ is the factor used to control the balance between these two distances and is chosen to be 1 for convenience. The process only calculates the distance between points and predicted values of the surrogate model on the preselected data set, resulting in a negligible calculation time. Then, we have $\gamma^{j+1}=\gamma^{j}\cup\{\widetilde{m}_{j+1}\}$. We can improve the DeepOnet model $\mathcal{F}_{\theta}$ by collecting the new training set $\gamma^{Q}$. Specifically, during the online training phase, we initialize the neural network parameter from the previously trained model $\mathcal{F}_{\theta}$. We expect a significant speedup in solving Eq. (5) with this initialization, which can be viewed as an example of transfer learning.
The advantages of this method are obvious. Specifically, the selected adaptive points will have varying features in the surrogate solution spaces and will be close to the anchor point $r^{t}$. This is more of an adversarial process; in order to guarantee local accuracy, we want the new points to be close to the $r^{t}$. In order to enhance generalization capacity, we also want them to incorporate as much data as the current surrogate did. The only remaining question is how to approximate the approximate posterior distribution $\widehat{\nu}_{t}$. To this end, traditional MCMC-based sampling methods and particle-based approaches[44, 45, 36, 46, 35] can be applied. In this paper, we have chosen to exclusively focus on the Unscented Kalman Inversion (UKI) method [38] in order to conceptually validate that our proposed adaptive operator learning framework. The details of the UKI algorithm will be presented in the following section. It is important to emphasize that UKI’s strategy relies on Gaussian approximations. While UKI may not provide highly accurate posterior approximations in many cases, such as non-Gaussian posteriors, we have still chosen to use UKI for sampling due to its computational efficiency. As a gradient-free, particle-based method, UKI requires only $12N_{m}+$ 1 full-order model evaluations per iteration and typically converges within $\mathcal{O}(10)$ iterations, making it computationally less expensive compared to MCMC-type methods. The numerical examples will demonstrate that even with the lower computational complexity of UKI, our newly designed framework still achieves significant improvements in computational efficiency. Additionally, our algorithm is easily integrated with MCMC or other particle-based methods[47, 48, 49, 50, 36, 37, 51, 52, 53], and when used with other samplers that require a larger number of samples, our framework yields even greater computational efficiency gains.
@@ 3.2 Unscented Kalman Inversion
Let $Y_{n}:=\{y_{1},y_{2},\cdots,y_{n}\}$ denote the observation set at time $n$. In order to approximate the conditional distribution $\nu_{n}$ of $m_{n}|Y_{n}$, the iterative algorithm starts from the prior $\nu_{0}$ and updates $\nu_{n}$ through the prediction and analysis steps: $\nu_{n}\rightarrow\widetilde{\nu}_{n+1}$, and then $\widetilde{\nu}_{n+1}\rightarrow\nu_{n+1}$, where $\widetilde{\nu}_{n+1}$ is the distribution of $m_{n+1}|Y_{n}$. In the prediction step, we assume that $\nu_{n}=\mathcal{N}(r_{n},C_{n})$, then under Eq. (12), $\widetilde{\nu}_{n+1}$ is also Gaussian with mean and covariance:
@@
We obtain the following UKI algorithm in Algorithm 1 by applying the aforementioned quadrature rules. UKI is a derivative-free algorithm that applies a Gaussian approximation theorem iteratively to transport a set of particles to approximate given distributions. As a result, it only needs $12N_{m}+$ 1 model evaluations per iteration, making it simple to implement and inexpensive to compute. However, for highly-nonlinear problems, UKI may encounter the intrinsic difficulty of using Gaussian distributions to approximate the posterior and different samplers can be applied.
@@ 3.4 Convergence analysis under the linear case
Under Assumptions 3.2-3.4, suppose $Range(\mathcal{G}^{T})=Range(\widehat{\mathcal{G}}^{T})=\mathbb{R}^{N_{m}}$ and $\Sigma_{\omega}\succ 0,\Sigma_{\eta}\succ 0$. Then the sequence $\widehat{r}_{\infty},\widehat{C}_{\infty}^{-1}$ in Eq.(18) obtained by using the surrogate model will converge to the $r_{\infty},C_{\infty}^{-1}$ in Eq.(17) and we have the following error estimate
@@ 4 Numerical experiments
$y_{obs}=y_{ref}+\mbox{max}\{|y_{ref}|\}\delta\xi,$ (27)
where $y_{ref}=\mathcal{G}(m_{ref})$ are the exact data, $\delta$ dictates the relative noise level and $\xi$ is a Gaussian random variable with zero mean and unit standard deviation. In UKI, the regularization parameters are $\alpha=0.5$ for noise levels 0.05 and 0.1 and $\alpha=1$ for noise levels 0.01. The starting vector for the UKI is chosen at random from $\mathcal{N}(0,I)$. The selection of other hyperparameters is based on [38]. For numerical examples, we set $\Omega=[0,1]^{2}$. The maximum number of UKI iterations per cycle is 20 for FEM-UKI and DeepOnet-UKI-Direct, and 10 for DeepOnet-UKI-Adaptive. Following that, we will use the greedy algorithm to select $Q=50$ adaptive samples for noise level 0.01 and $Q=20$ adaptive samples for noise levels 0.05 and 0.1 from $K=2000$ samples.
@@
to demonstrate that our adaptive framework can actually reduce the local model error. Moreover, we also calculate the data fitting error via Eq.(3) using the true model during the UKI iteration process.
@@ 4.1 Example 1: Darcy flow
[…]
Figure 4: The ground truth for in-distribution data(IDD) and out-of-distribution data (OOD) from above to below. Left: the true permeability fields $\mathbf{m}(\mathbf{x})$. Right: the pressure fields $u(\mathbf{x})$ and the corresponding $36$ equidistant observations with noise level 0.01.
@@
$\mathbf{m}(\mathbf{x})=\sum_{k\in\mathbb{Z}^{+}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\zeta_{k}}\sqrt{\lambda_{k}}\psi_{k}(\mathbf{x}),$ (33)
where $\lambda_{k}$ and $\psi_{k}$ are the eigenvalues and eigenfunctions, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\zeta_{k}}\sim\mathcal{N}(0,1)$ are independent random variables. In practice, we truncate the sum (33) to $n_{d}$ terms, based on the largest $n_{d}$ eigenvalues, and hence ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\zeta}\in\mathbb{R}^{n_{d}}$. The forward problem is solved by FEM method on a $70\times 70$ grid.
We will create the observation data for the inverse problem using the in-distribution data (IDD) and out-of-distribution data (OOD), respectively, as shown in Fig.4. The IDD field $\mathbf{m}_{ref}(\mathbf{x})$ is calculated using Eq.(33) with $n_{d}=256$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\zeta_{k}}\sim\mathcal{N}(0,1)$. The OOD field $\mathbf{m}_{ref}(\mathbf{x})$ is generated for convenience by sampling ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\zeta_{k}}\sim\mathcal{U}[-20,20],k=1,\ldots,256$. To avoid the inverse crime, we will try to inverse the first $N_{m}=128$ KL modes using these observation data.
@@
[…]
Figure 5: The data fitting error (left), model error (middle) and relative inversion error (right) for Example 1. Above: IDD case. Below: OOD case.
[…]
Figure 6: IDD case. Above: the estimated permeability field obtained by different methods. Below: the absolute errors with respect to the true ones.
[…]
Figure 7: OOD case. Above: the estimated permeability field obtained by different methods. Below: the absolute errors with respect to the true ones.
[…]
Figure 8: The error box plot of $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$ with different numbers of points $Q$(above) and noise levels (below). Left: IDD case. Right: OOD case.
[…]
Figure 9: The mean total number of forward evaluations for FEM-UKI and DeepOnet-UKI- Adaptive respectively. Left: IDD case and Right: OOD case.
[…]
Figure 10: Left: the average computational time and average number of forward evaluations each iteration for IDD case. Right: the average computational time and average number of forward evaluations each iteration for IDD case. Current noise level is 0.01.
To test the effect of the number $Q$ of adaptive samples used in each refinement, we repeat the experiment ten times for each $Q\in[20,50,100,150]$. The error box of $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$ is plotted Fig.8. It is evident from the IDD data that the relative inversion error does not decrease significantly with increasing data. This suggests that a small set of adaptive samples– roughly 50 – can meet the requirements for accuracy and efficiency. On the other hand, the relative inversion error for OOD data steadily drops with increasing dataset size. In order to examine the effects of varying noise levels, we are going to perform the experiment with three different noise levels (0.01, 0.05, and 0.1) and then repeat it with ten different UKI initial values. The numerical results are also shown in Fig.8. We can clearly observe that the relative inversion error gradually drops as noise levels rise, suggesting that higher noise levels are less sensitive to model errors. Consequently, our framework performs better in real-world applications with higher noise levels.
To test the computational efficiency of our adaptive framework, we plot the mean total online forward evaluations in Figure 9. Our method incurs a significantly lower cost compared to conventional numerical methods, even when adaptive refinement is applied. Specifically, our approach requires a maximum of only 50 samples for model refinement, whereas FEM-UKI demands 5,140 forward evaluations. To further evaluate the computational efficiency of our adaptive framework, we present the average computational CPU times and the number of forward evaluations per valid iteration in Figure 10. The results show that, apart from the initial offline cost for generating data points and training, the iterative computational efficiency of our adaptive framework surpasses traditional FEM solvers for both IDD and OOD cases. Given that the online fine-tuning process is highly efficient, with retraining taking just a few seconds, the cost of these forward simulations becomes negligible. Moreover, even when considering the offline cost, the total number of forward evaluations required by our framework is substantially lower than that of FEM-UKI, indicating that our scheme offers superior efficiency, particularly when the underlying PDE is expensive to solve.
@@ 4.2 Example 2: The heat source inversion problem
The objective is to identify the heat source $f$ from noisy measurements. To illustrate our method more clearly, we divide this inversion task into two cases. In both cases, the surrogate is the DeepOnet model with same architectures. The FEM method is used to solve the forward problem on a $70\times 70$ grid and the resulting differential equations are integrated using the implicit-Euler scheme.
Case I: In this case, we consider a 2D heat source inversion problem which is adapted from[11]. Take $f=\frac{s}{2\pi\sigma^{2}}\exp\left(-\frac{\left|\boldsymbol{\chi}-\mathbf{x}\right|^{2}}{2\sigma^{2}}\right)[1-H(t-T)]$ with zero initial condition and zero Neumann boundary condition, where $H$ is the heavy-side function. Take $s=5,\sigma=0.1,T=0.05$. The inverse problem is to infer the location by giving the observations at $t=0.05$ and $t=0.15$. In this paper, we take the ground truth to be $\mathbf{x}_{true}=[0.2,0.2]$. A uniform $3\times 3$ sensor network is used to collect noisy point-wise observations from the PDE field solution. At each sensor location, two measurements are taken at $t=0.05$ and $t=0.15$, for a total of 18 measurements.
To replace the forward model, DeepOnet is trained on $[0.5,1]\times[0.5,1]$ with 500 uniformly distributed samples. We use UKI with our adaptive scheme to run the experiment with initial value $[0.6,0.6]$. The numerical results are shown in Fig.11. The DeepOnet-UKI-Direct method can provide only a rough estimate. While with our adaptive refinement, the DeepOnet-UKI-Adaptive can nearly achieve the same accuracy as FEM-UKI. That is because the adaptive sampling can select import samples in the high density area of the approximate posterior and the surrogate is refined to reduce the local model error. This phenomenon can be further observed in Fig.12, which plots the inversion trajectories and samples distribution for our adaptive method. We can clearly see that with refinements, the surrogate gradually become accurate in the high density area of the true posterior, which can modify the inversion trajectory and thus lead to more accurate results compared to DeepOnet-UKI-Direct.
@@
[…]
Figure 11: The fitting error, model error and relative inversion error from left to right respectively.
[…]
Figure 12: The inversion trajectories of the three methods and the samples distribution of DeepOnet-UKI-Adaptive.
Case II: In this case, the heat source field is considered with the formula $e^{-t}\mathbf{m}(\mathbf{x})$ with zero Dirichlet boundary condition and initial condition ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u_{0}(x,y)}=100\sin(x)\sin(y)$. Conversely, the inverse problem involves using noisy measurements of $u(\mathbf{x},1)$ to determine the true spatial source field $\mathbf{m}(\mathbf{x})$. We assume that the Gaussian random field defined in Eq.(33) is the prior of $\mathbf{m}(\mathbf{x})$.
We assume that the ground truth $\mathbf{m}_{ref}(\mathbf{x})$ has an analytical solution in this example to increase the dimension of the problem, i.e.,
@@
Figure 13: Example 2. Left: the true source field. Right: the temperature field $u(\mathbf{x},1)$ and the corresponding $36$ equidistant observations with noise level 0.01.
To test the effectiveness of our framework, we first run the experiment using the original pre-trained model directly running UKI, i.e. DeepOnet-UKI-Direct. We plot the local model error and the relative inversion error in the middle and right displays of Fig.14, respectively. As expected, the local model error will increase significantly, and the pre-trained model will eventually fail to predict the result. Because of the growing model error, the relative inversion error follows accordingly, increasing dramatically and providing a totally inaccurate final estimate. Nonetheless, we may refine our model by constructing adaptive samples based on this estimate. The procedures are similar to Example 1. During the exploration stage, we will run UKI for 10 steps and select the anchor point with the smallest data fitting error as the initial value for the next run of UKI. To reduce model error, we will generate adaptive samples near the anchor point and refine the surrogate. As shown in the middle of Fig.14, the local model error significantly decreases after refinement. As a result, our method produces relative inversion errors that are significantly reduced after refinement and comparable to those of FEM-UKI. Finally, after several refinements, the entire procedure is terminated based on the stop criteria, with the black star indicating the final point that we would accept for the inversion process using our DeepOnet-UKI-Adaptive algorithm. To demonstrate the effectiveness of our method, we plot the inversion results in Figs.15 and 16. We can see that the final numerical results produced by DeepOnet-UKI-Adaptive and FEM-UKI are very similar and do not differ significantly. This suggests that our method can handle OOD data in closed form.
@@
Figure 14: The inversion results obtained by the three methods for the heat source case. Left: the data fitting error. Middle: the model error. Right: the relative inversion errors.
[…]
Figure 15: The inversion results for the heat source case. Above: the estimated source fields obtained by three different methods. Below: the absolute errors with respect to the true one.
[…]
Figure 16: The inversion results for the heat source case. Above: the approximate states evaluated at the final estimated source fields obtained by different methods. Below: the absolute errors between the approximate state and the true state.
[…]
Figure 17: The inversion results for the heat source case. Left: the error box plot of $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$. Middle: the mean total number of forward evaluations. Right: the average computational time.
In order to provide additional evidence of the efficacy of our approach, we figure out to perform the experiment for UKI at varying noise levels. In addition, we will repeat the experiment ten times with different initial values for each noise level. Following that, we will compare the number of forward evaluations and the relative inversion errors for every approach. We plot the difference of the relative inversion errors $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$ in the left display of Fig.17. It is clear that when noise levels increase, DeepOnet-UKI-Adaptive performs often better than FEM-UKI. This implies that our method can achieve higher accuracy than traditional solvers. For the reasons mentioned below, the computational cost of the new method can also be extremely small. First of all, it is much faster to fine-tune the original pre-trained surrogate model than it is to solve PDEs. Specially, we only need a maximum of 50 online forward evaluations in this example to retrain the network, which drastically lowers computational costs. We are able to clearly see that DeepOnet-UKI-Adaptive has a substantially smaller total number of forward evaluations than FEM-UKI, as the middle display of Fig.17 illustrates. Secondly, the entire process is automatically stopped by applying the stop criterion. We can start with the initial model that has been trained offline and fine-tune it multiple times for a given inversion task. As a result, our framework can achieve an accuracy level comparable to traditional FEM solvers, but at a significant reduction in computational cost for such problems. The CPU computation times are plotted in the right display of Fig.17. It is clear that our adaptive framework greatly accelerates the inversion process, which is nearly 10 times faster than FEM-UKI with nearly same accuracy.